# Anchor-based Multi-view Subspace Clustering with Hierarchical Feature Descent

### v1

@@ Abstract
Multi-view clustering has attracted growing attention owing to its capabilities of aggregating information from various sources and its promising horizons in public affairs. Up till now, many advanced approaches have been proposed in recent literature. However, there are several ongoing difficulties to be tackled. One common dilemma occurs while attempting to align the features of different views. We dig out as well as deploy the dependency amongst views through hierarchical feature descent, which leads to a common latent space( STAGE 1). This latent space, for the first time of its kind, is regarded as a “resemblance space”, as it reveals certain correlations and dependencies of different views. To be exact, the one-hot encoding of a category can also be referred to as a resemblance space in its terminal phase. Moreover, due to the intrinsic fact that most of the existing multi-view clustering algorithms stem from k-means clustering and spectral clustering, this results in cubic time complexity w.r.t. the number of the objects. However, we propose Anchor-based Multi-view Subspace Clustering with Hierarchical Feature Descent(MVSC-HFD) to further reduce the computing complexity to linear time cost through a unified sampling strategy in resemblance space( STAGE 2), followed by subspace clustering to learn the representation collectively( STAGE 3). Extensive experimental results on public benchmark datasets demonstrate that our proposed model consistently outperforms the state-of-the-art techniques.
@@
[…]
organization=School of Computer, National University of Defense Technology,addressline=ouqiyuan14, wangsiwei13, zhangpei, enzhu@nudt.edu.cn, sihangjoe@gmail.com, city=Changsha, postcode=410073, state=Hunan, country=China
{graphicalabstract}{highlights}
MVSC-HFD has high adaptability to variant dimensions.
Reduce multiple views of different dimensions hierarchically to a unified space.
Uniformly learn anchors in embedding space and construct bipartite graphs.
@@
Variables are collectively learned through optimization with guaranteed convergence.
@@ 1 Introduction
With the advance of information technology, many practical datasets are collected from different channels or provided in different modalitiesCai et al. (2023); Liang et al. (2022); Yang et al. (2022b); Xu et al. (2013). As a result, multi-view learning has gained increasing attention due to its ability in handling and learning from various source information. Among those learning datasets, the majorities are unlabeled with only features of different views attached to them. Under the obligation to discover the inherent pattern from these cluttered datasets and in aspiration for bringing to light what used to be hidden from human conception, many clustering algorithms are deployed in practice. One common task that multi-view clustering algorithms confront is that in some datasets, different views vary tremendously in dimensions, which will adversely affect the extraction and expression of common representations and data patterns revealed by features from each view.
Projecting different views into a common space is a long-term academic issue, e.g., Canonical Correlation Analysis (CCA)Hotelling (1992) and its variants Akaho (2006); Andrew et al. (2013) aim to align a pair of views with linear correlation through complex nonlinear transformation. This technique is restricted to the simple case where only two views are provided. In multi-view circumstances, Wang et al. (2020) assumes that the high dimensional data emanate from noisy redundancies, which can be eliminated by minimizing $l_{2,1}$ norm of the representation. It is a typical phenomenon that, the lower the dimensions of data points are reduced to, the more the information is lost Qaqish et al. (2017); Wei et al. (2023). On the contrary, the higher dimensions we preserve to learn from, the better the resulting representation captures the underlying semantic essence of the data points, although with typically higher computational cost.
@@
Subspace clustering directly bypasses the curse of dimensionalityBellman (1954) through expressing each data point as a linear combination of other points and minimizing the reconstruction loss to attain the combination coefficientLiu et al. (2013); Elhamifar and Vidal (2013). Recently, subspace clustering has been extended to multi-view circumstancesGao et al. (2015a); Cao et al. (2015); Yan et al. (2023); Chen et al. (2022); Zhao et al. (2023), where each view owns a distinct coefficient matrix and a regularization term is used to obtain the joint coefficient matrix, which is then expressed as similarity matrix for downstream spectral clustering and k-means clustering tasks. It is obvious that, as the self-representation matrix has the size of $n\times n$, which is the same as the scale of data points, extremely high computational and spatial cost are imposed on the optimization procedure and the clustering task, which would bring $\mathcal{O}(n^{3})$ time cost and at least $\mathcal{O}(n^{2})$ storage cost, respectively. As a typical paradigm for reducing the computational and spatial cost, the anchor scheme is introduced to various clustering modelsChen and Cai (2011); Wang et al. (2016); Kang et al. (2020); Zhou et al. (2021); Shu et al. (2022); Yang et al. (2023). It is a common practice to carefully choose (mainly adopt k-means clustering as a preliminary scheme) or randomly sample from the original dataset. The above formula bears several drawbacks: a) Results highly rely on the initialization as the anchors are predetermined. Although a preliminary $k$-means clustering on data features can make a favorable selection, the $\mathcal{O}(n^{3})$ pretreat cost greatly reduces their practicability; b) In multi-view learning background, the fundamental issue that underlies most of the frameworks is that both the anchors and the graph reconstruction matrices are view-specific, which neglects the alliance amongst different views; c) For final spectral clustering step, the fusion step is required to build the global graph based on view-specific graphs. In this circumstance, the fusion method will also extensively affect the clustering result.
To address the aforementioned limitations, we propose a novel approach named Anchor-based Multi-view Subspace Clustering with Hierarchical Feature Descent(MVSC-HFD). Specifically, in our method, we adopt three mechanisms to extract the global graph reconstruction matrix. First, the input data points of different views are reduced to the same dimensional space through hierarchical dimensional projection matrices. Second, we assume that for all views, data points portray/sketch the same structure in this unified dimension, and can be jointly expressed by common subspace, thus reuniting the separate clustering of distinct views into a collectively learned framework. Thirdly, to negotiate between feature preservation and efficiency, we adopt an anchor graph with carefully designed dimensions and the number of anchor points. The advantages of our proposed algorithm are listed below:
@@
    Through uniformly learned anchors in embedding space, MVSC-HFD constructs the bipartite graph between landmarks and original data samples. Our algorithm has low spatial and time consumption which are linear to data size.
@@
    By means of hierarchically adjusting the dimension of data representation, the dimensional reduction process reduces redundancy in features from multiple views. The common low-dimensional embedding space maintains the dependency whereas handling the discrepancy among different views.
@@
    During optimization of the clustering process, the optimal weights of each view are automatically learned with no additional parameters nor pre-specified knowledge needed for adjustment.
@@ 2 Related Work
$d_{v}$ Dimensionality of the $v$-th view.
$l_{i}$ Dimensionality of the $i$-th layer.
@@ 2.1 Non-Negative Matrix Factorization
Non-negative matrix factorization (NMF or NNMF) is a group of algorithms that factorize a matrix $V$ into (usually) two matrices $F$ and $H$, with the uniform constraint that all these three matrices have the property of non-negativity, which makes the formulation easy to solve whereas endow the practicality of the algorithms in real-world applications such as processing of natural language, audio-visual signal spectrograms or missing data imputation.
${V}\approx{FH},$ (1)
where ${V}\in\mathbb{R}^{d\times n}$ is the data matrix, each column represents a vector of dimension $d$ and there are $n$ numbers of such vector. ${F}\in\mathbb{R}^{d\times k}$ and $k$ is the number of base vectors. ${H}\in\mathbb{R}^{k\times n}$ is the representation mapped through the mapping matrix ${F}$. While keeping the information unchanged as much as possible, NMF aims to simplify high dimensional data $V$ to low dimensional mode $H$, in order to estimate the essential structure which $V$ represents. Original NMF and Deep-NMF are illustrated in Fig. 1.
@@ 2.3 Multi-view Subspace Clustering
where the regularization term $\|\cdot\|_{\dagger}$ endows view-specific anchor graphs $\{\mathbf{S}_{i}\}_{i=1}^{p}$ with different properties. $\mathbf{S}$ is the self-representation matrix optimized by minimizing the term $\Omega(\mathbf{S},\mathbf{S}_{i})$, which represents categories of many fusion techniques. The clustering process of existing methods for multi-view subspace clustering is in two stages: - Step 1: Learning the subspace representation and integrating multi-view information at the level of similarity matrix or self-expression matrix by fusing the graphs or self-expression matrices of multiple views (reconstruction matrix) to form a shared similarity matrix Li et al. (2015); Wang et al. (2015, 2018). For example, Guo (2013) learned shared sparsity by performing Matrix Factorization (MF) of the subspace representations. Unlike obtaining shared self-expressions or shared graphs directly, to better exploit the complementary information and diversity between different views, the work Cao et al. (2015); Huang et al. (2021) and Xia et al. (2014) were performed by Hilbert-Schmidt independence criterion (HSIC) and Markov chain, respectively. Wang et al.Wang et al. (2017) introduced an exclusion constraint to learn different subspace representations in each view, which force multiple subspace representations to be sufficiently different from each other, and then added them together directly or adaptively to obtain a uniform graph or self-expression matrix. - In the second step: Clustering, the obtained uniform graph representation is fed into an off-the-shelf clustering model (generally using spectral clustering) to obtain the final clustering results.
Although satisfying clustering performances have been achieved by existing multi-view subspace learning algorithms in many real world applications, the large computational consumption and storage cost by optimizing the $n\times n$ reconstruction matrix $\mathbf{S}$ prohibits the scalability of these methods on large-scale background.
@@ 3.1 Motivation
In large-scale data circumstances (either with large data size or with high dimensionality), the time and space overhead of traditional learning algorithms would be overwhelmingly hugeShakhnarovich (2005). For high dimensional datasets, it is of much redundancy to directly learn representations from original data samples. Although subspace clustering provides an alternative approach to circumvent high-dimensional feature learning, it still spends high computational and storage costs to construct global similarity graphs between each two elements. And it is obvious that a large proportion of the data does not contribute significantly to the clustering structure. Anchor graph theoryWang et al. (2016); Yang et al. (2022c) proposes that a good clustering structure can be acquired with only a fraction of the sample points by correlating these sample points with the full sets of data. These similarity matrices are also termed as anchor graph or bipartite graph. Nevertheless, existing anchor-based subspace learning methods often use delicately chosen landmarks beforehand, hence the clustering result highly relies upon these pre-specified knowledge, which restricts the scalability of the algorithm. We propose an adaptable clustering algorithm to tackle not only high dimensional data, but also large-scale samples by first projecting the original data to a low dimensional embedding space, where a set of anchors are learned uniformly to construct the anchor graph. We demonstrate the overall procedure together with traditional methods in fig. 2.
@@
Compared with existing methods, our algorithm has the following advantages:
  • 1.
    High adaptability to variant dimensions. Our algorithm can reduce multiple views of different dimensions hierarchically at different rates into a uniform dimensional space for clustering.
  • 2.
    Anchor learnability and view-consistent anchor graphs. We use anchors, which are learned in a uniform low-dimensional space, to further obtain cross-view consistent bipartite graph with the full set of sample points without additional graph fusion.
  • 3.
    Variables integrity and parameter learnability. Aforementioned variables including projection matrix, anchors and anchor graphs are collectively learned throughout the optimization procedure with guaranteed convergence.
@@ 3.2 The Proposed Formulation
where $\mathbf{X}^{(v)}\in\mathbb{R}^{d_{v}\times n}$ is the original data matrix of the $v$-th view, with $d_{v}$ denoting the dimension and $n$ denoting the number of data points. $\mathbf{A}\in\mathbb{R}^{k\times m}$ is the unified anchor points in which $k$ represents the dimensionality and $m$ stands for the number of the designated anchor points. The hierarchical dimensional projection matrix $\mathbf{W}^{(v)}_{i}\in\mathbb{R}^{l_{i-1}\times l_{i}}$ ($l_{0}=d_{v}$, $l_{\delta}=k$, in particular) is the projection matrix that performs dimensional reduction. The learnable vector of parameters $\boldsymbol{\alpha}=[\alpha_{1},\alpha_{2},\cdots,\alpha_{p}]^{\top}$ represents the weights adherent to different views. And the linear constraint of $\boldsymbol{\alpha}$ corresponds to the quadratic optimization target, which is a convex optimization problem owing to the inherently non-negative nature of Frobenius norm. $\mathbf{Z}\in\mathbb{R}^{m\times n}$ represents the consensus bipartite graph amongst different views, which constructs the pairwise relationship between original data samples and delicately learned landmarks. Traditional subspace clustering algorithms utilize the learned bipartite graph to construct the full similarity graph $\mathbf{S}$ using $\mathbf{S}=\mathbf{\hat{Z}^{\top}\hat{Z}}$ where $\mathbf{\hat{Z}}=\mathbf{Z}\Sigma^{-1/2}$ and $\Sigma$ denotes a diagonal matrix with $\Sigma_{jj}=$ $\sum_{i=1}^{n}\mathbf{Z}_{ij}$. Afterwards, they perform spectral clustering with the acquired similarity matrix to obtain final clustering results. These give rise to $\mathcal{O}(n^{3})$ computational consumption and $\mathcal{O}(n^{2})$ spatial cost. We adopt a different approach by proving an equivalence in right singular matrix of bipartite graph $\mathbf{Z}$ and eigenvectors of the similarity matrix $\mathbf{S}$:
$\begin{split}\mathbf{S}=\mathbf{\hat{Z}^{\top}\hat{Z}}&=(\left(U\Lambda V^{\top}\right)\Sigma^{-1/2})^{\top}\left(U\Lambda V^{\top}\Sigma^{-1/2}\right)\\ &=V\Lambda\Sigma^{-1/2}\left(U^{\top}U\right)\Sigma^{-1/2}\Lambda V^{\top}\\ &=V\Lambda^{2}\Sigma^{-1}V^{\top},\end{split}$ (7)
therefore, we only need to compute the right singular matrix of $\mathbf{Z}\in\mathbb{R}^{m\times n}$ for further clustering task (lite $k$-means for example), and this costs $\mathcal{O}(nm^{2})$.

### v2

@@ Abstract
Multi-view clustering has attracted growing attention owing to its capabilities of aggregating information from various sources and its promising horizons in public affairs. Up till now, many advanced approaches have been proposed in recent literature. However, there are several ongoing difficulties to be tackled. One common dilemma occurs while attempting to align the features of different views. Moreover, due to the fact that many existing multi-view clustering algorithms stem from spectral clustering, this results to cubic time complexity w.r.t. the number of dataset. However, we propose Anchor-based Multi-view Subspace Clustering with Hierarchical Feature Descent(MVSC-HFD) to tackle the discrepancy among views through hierarchical feature descent and project to a common subspace( STAGE 1), which reveals dependency of different views. We further reduce the computational complexity to linear time cost through a unified sampling strategy in the common subspace( STAGE 2), followed by anchor-based subspace clustering to learn the bipartite graph collectively( STAGE 3). Extensive experimental results on public benchmark datasets demonstrate that our proposed model consistently outperforms the state-of-the-art techniques. Our code is publicly available at https://github.com/QiyuanOu/MVSC-HFD/tree/main.
@@
MVSC-HFD has high adaptability to various dimensions of different views.
Tackle the discrepency among different views.
Reduce multiple views of different dimensions hierarchically to a common subspace.
Uniformly learn anchors in common subspace and construct bipartite graphs.
@@ 1 Introduction
With the advance of information technology, many practical datasets are collected from different channels or provided in different modalitiesCai et al. (2023); Liang et al. (2022); Yang et al. (2022b); Xu et al. (2013). As a result, multi-view learning has gained increasing attention due to its ability in handling and learning from various source information. Among those learning datasets, the majorities are unlabeled with only features of different views attached to them. Many clustering algorithms have been used in practice in order to discover inherent patterns from these cluttered datasets and also to reveal things that were previously unknown to humans. One common task that multi-view clustering algorithms confront is that in some datasets, different views vary tremendously in dimensions, which will adversely affect the extraction and expression of common representations and data patterns revealed by features from each viewYan et al. (2023a, c); Tang et al. (2023b); Yan et al. (2022); Tang et al. (2023c, 2022, a). Projecting different views into a common space is a long-term academic issue, e.g., Canonical Correlation Analysis (CCA)Hotelling (1992) and its variants Akaho (2006); Andrew et al. (2013) aim to align a pair of views with linear correlation through complex nonlinear transformation. This technique is restricted to the simple case where only two views are provided. In multi-view circumstances, Wang et al. (2020) assumes that the high dimensional data emanate from noisy redundancies, which can be eliminated by minimizing $l_{2,1}$ norm of the representation. It is a typical phenomenon that, the lower the dimensions of data points are reduced to, the more the information is lost Qaqish et al. (2017); Wei et al. (2023). On the contrary, the higher dimensions we preserve to learn from, the better the resulting representation captures the underlying semantic essence of the data points, although with typically higher computational cost. Subspace clustering directly bypasses the curse of dimensionalityBellman (1954) through expressing each data point as a linear combination of other points and minimizing the reconstruction loss to attain the combination coefficientLiu et al. (2013); Elhamifar and Vidal (2013). Recently, subspace clustering has been extended to multi-view circumstancesGao et al. (2015a); Cao et al. (2015); Yan et al. (2023b); Chen et al. (2022); Zhao et al. (2023), where each view owns a distinct coefficient matrix and a regularization term is used to obtain the joint coefficient matrix, which is then expressed as similarity matrix for downstream spectral clustering and k-means clustering tasks. It is obvious that, as the self-representation matrix has the size of $n\times n$, which is the same as the scale of data points, extremely high computational and spatial cost are imposed on the optimization procedure and the clustering task, which would bring $\mathcal{O}(n^{3})$ time cost and at least $\mathcal{O}(n^{2})$ storage cost, respectively. As a typical paradigm for reducing the computational and spatial cost, the anchor scheme is introduced to various clustering modelsChen and Cai (2011); Wang et al. (2016); Kang et al. (2020); Zhou et al. (2021); Shu et al. (2022); Yang et al. (2023). It is a common practice to carefully choose (mainly adopt k-means clustering as a preliminary scheme) or randomly sample from the original dataset. The above formula bears several drawbacks: a) Results highly rely on the initialization as the anchors are predetermined. Although a preliminary $k$-means clustering on data features can make a favorable selection, the pre-selected anchor points are fixed in the same dimensional space as the original data, so the practicability can be greatly reduced when the nodes are not linearly separable; b) In multi-view learning background, the fundamental problem with most of the frameworks is that both the anchors and the graph reconstruction matrices are view-specific, which ignores the alliance between different views; c) For final spectral clustering step, the fusion step is required to build the global graph based on view-specific graphs. In this circumstance, the fusion method will also extensively affect the clustering result. d) Specifically for Wang et al. (2021); Li et al. (2020), which efficiently learn anchors using subspace clustering method, assume that cross-view consistent bipartite graph exists. Accordingly, they deploy dependency among views without tackling the discrepancy of dimensions and features adequately, thus inhibiting a better performance. To address the aforementioned limitations, we propose a novel approach named Anchor-based Multi-view Subspace Clustering with Hierarchical Feature Descent(MVSC-HFD). Specifically, in our method, we adopt three mechanisms to extract the bipartite graph reconstruction matrix. First, the input data points of different views are reduced to the same dimensional space through hierarchical dimensional projection matrices. Second, we assume that for all views, data points sketch the same structure in this unified dimension, and can be jointly expressed by common subspace, thus reuniting the separate clustering of distinct views into a collectively learned framework. Thirdly, to negotiate between feature preservation and efficiency, we adopt an anchor graph with carefully designed dimensions and the number of anchor points. Compared with existing methods, our algorithm has the following advantages:
@@
    High adaptability to various dimensions. Our algorithm can cater to different dimensions of multiple views.
@@
    Reduce the discrepancy and reveal dependency of different views. Our method can tackle discrepancy among multiple views by hierarchically dimensional descent at different rates into a common subspace for clustering, which increases the effectiveness of the anchors-based clustering scheme.
@@
    Model integrity and performance gains. Aforementioned variables including weight vector, projection matrix, anchor matrix and bipartite graph are collectively learned throughout the optimization procedure with guaranteed convergence. The extensive experiments show the validity of our proposed method from both accuracy and efficiency.
@@ 2 Related Work
$d_{v}$ Dimension of the $v$-th view.
$l_{i}$ Dimension of the $i$-th layer.
@@ 2.1 Non-Negative Matrix Factorization
Non-negative matrix factorization (NMF or NNMF) is a group of algorithms that factorize a matrix $\mathbf{V}$ into (usually) two matrices $\mathbf{F}$ and $\mathbf{H}$, with the uniform constraint that all these three matrices have the property of non-negativity, which makes the formulation easy to solve whereas endow the practicality of the algorithms in real-world applications such as processing of natural language, audio-visual signal spectrograms or missing data imputation.
$\mathbf{V}\approx\mathbf{FH},$ (1)
where $\mathbf{V}\in\mathbb{R}^{d\times n}$ is the data matrix, each column represents a vector of dimension $d$ and there are $n$ numbers of such vector. $\mathbf{F}\in\mathbb{R}^{d\times k}$ and $k$ is the number of base vectors. $\mathbf{H}\in\mathbb{R}^{k\times n}$ is the representation mapped through the mapping matrix $\mathbf{F}$. While keeping the information unchanged as much as possible, NMF aims to simplify high dimensional data $\mathbf{V}$ to low dimensional mode $\mathbf{H}$, in order to estimate the essential structure which $\mathbf{V}$ represents. Original NMF and Deep-NMF are illustrated in Fig. 1.
@@ 2.3 Multi-view Subspace Clustering
where the regularization term $\|\cdot\|_{\dagger}$ endows view-specific anchor graphs $\{\mathbf{S}_{i}\}_{i=1}^{p}$ with different properties. $\mathbf{S}$ is the self-representation matrix optimized by minimizing the term $\Omega(\mathbf{S},\mathbf{S}_{i})$, which represents categories of many fusion techniques. The clustering process of existing methods for multi-view subspace clustering is in two stages: - Step 1: Learning the subspace representation and integrating multi-view information at the level of similarity matrix or self-expression matrix by fusing the graphs or self-expression matrices of multiple views (reconstruction matrix) to form a shared similarity matrix Li et al. (2015); Wang et al. (2015, 2018). For example, Guo (2013) learned shared sparsity by performing Matrix Factorization (MF) of the subspace representations. Unlike obtaining shared self-expressions or shared graphs directly, to better exploit the complementary information and diversity between different views, the work Cao et al. (2015); Huang et al. (2021) and Xia et al. (2014) were performed by Hilbert-Schmidt independence criterion (HSIC) and Markov chain, respectively. Wang et al.Wang et al. (2017) introduced an exclusion constraint to learn different subspace representations in each view, which force multiple subspace representations to be sufficiently different from each other, and then added them together directly or adaptively to obtain a uniform graph or self-expression matrix. - In the second step: Clustering, the obtained uniform graph representation is fed into an off-the-shelf clustering model (generally using spectral clustering) to obtain the final clustering results. Although satisfying clustering performances have been achieved by existing multi-view subspace learning algorithms in many real world applications, the large computational consumption and storage cost by optimizing the $n\times n$ reconstruction matrix $\mathbf{S}$ prohibits the scalability of these methods on large-scale background.
@@ 3.1 Motivation
In large-scale data circumstances (either with large data size or with high dimension), the time and space overhead of traditional learning algorithms would be overwhelmingly hugeShakhnarovich (2005). For high dimensional datasets, it is of much redundancy to directly learn representations from original data samples. Although subspace clustering provides an alternative approach to circumvent high-dimensional feature learning, it still spends high computational and storage costs to construct global similarity graphs between each two elements. And it is obvious that a large proportion of the data does not contribute significantly to the clustering structure. Anchor graph theoryWang et al. (2016); Yang et al. (2022c) proposes that a good clustering structure can be acquired with only a fraction of the sample points by correlating these sample points with the full sets of data. These similarity matrices are also termed as anchor graph or bipartite graph. Nevertheless, existing anchor-based subspace learning methods often use delicately chosen landmarks beforehand, hence the clustering result highly relies upon these pre-specified knowledge, which restricts the scalability of the algorithm. We propose an adaptable clustering algorithm to tackle not only high dimensional data, but also large-scale samples by hierarchically projecting the original data to a low dimensional common subspace, where a set of anchors are learned uniformly to construct the bipartite graph. We demonstrate the overall procedure together with traditional methods in fig. 2.
@@ 3.2 The Proposed Formulation
where $\mathbf{X}^{(v)}\in\mathbb{R}^{d_{v}\times n}$ is the original data matrix of the $v$-th view, with $d_{v}$ denoting the dimension and $n$ denoting the number of data points. The hierarchical dimensional projection matrix $\mathbf{W}^{(v)}_{i}\in\mathbb{R}^{l_{i-1}\times l_{i}}$ ($l_{0}=d_{v}$, $l_{\delta}=k$, in particular) is the projection matrix that performs dimensional reduction. Specifically, data in different views often have different dimensions. So usually $l_{i}$ are not the same for $\mathbf{W}^{(v)}_{i}$ of different views, which means different dimensional reduction process. Distinctive $\mathbf{W}^{(v)}_{i}$ represents the discrepancy of different views and the serial multiplier $\mathbf{W}^{(v)}_{1}\mathbf{W}^{(v)}_{2}\cdots\mathbf{W}^{(v)}_{\delta}$ aims to reduce the discrepancy. $\mathbf{A}\in\mathbb{R}^{k\times m}$ is the unified anchor points in which $k$ represents the dimension and $m$ stands for the number of the anchor points. This anchor matrix $\mathbf{A}$ is shared among different views and reveals the dependency. It represents the common features extracted through dimensional projection matrix multiplier $\mathbf{W}^{(v)}_{1}\mathbf{W}^{(v)}_{2}\cdots\mathbf{W}^{(v)}_{\delta}$. The learnable vector of parameters $\boldsymbol{\alpha}=[\alpha_{1},\alpha_{2},\cdots,\alpha_{p}]^{\top}$ represents the weights adherent to different views. And the linear constraint of $\boldsymbol{\alpha}$ corresponds to the quadratic optimization target, which is a convex optimization problem owing to the inherently non-negative nature of Frobenius norm. $\mathbf{Z}\in\mathbb{R}^{m\times n}$ represents the consensus bipartite graph amongst different views, which constructs the pairwise relationship between original data samples and delicately learned landmarks. Traditional subspace clustering algorithms utilize the learned bipartite graph to construct the full similarity graph $\mathbf{S}$ using $\mathbf{S}=\mathbf{\hat{Z}^{\top}\hat{Z}}$ where $\mathbf{\hat{Z}}=\mathbf{Z}\mathbf{\Sigma}^{-1/2}$ and $\mathbf{\Sigma}$ denotes a diagonal matrix with $\mathbf{\Sigma}_{jj}=$ $\sum_{i=1}^{n}\mathbf{Z}_{ij}$. Afterwards, they perform spectral clustering with the acquired similarity matrix to obtain final clustering results. These give rise to $\mathcal{O}(n^{3})$ computational consumption and $\mathcal{O}(n^{2})$ spatial cost. We adopt a different approach by proving an equivalence in right singular matrix of bipartite graph $\mathbf{Z}$ and eigenvectors of the similarity matrix $\mathbf{S}$:
$\begin{split}\mathbf{S}=\mathbf{\hat{Z}^{\top}\hat{Z}}&=(\left(\mathbf{U}_{\mathbf{Z}}\mathbf{\Lambda}_{\mathbf{Z}}\mathbf{V}_{\mathbf{Z}}^{\top}\right)\mathbf{\Sigma}^{-1/2})^{\top}\left(\mathbf{U}_{\mathbf{Z}}\mathbf{\Lambda}_{\mathbf{Z}}\mathbf{V}_{\mathbf{Z}}^{\top}\mathbf{\Sigma}^{-1/2}\right)\\ &=\mathbf{V}_{\mathbf{Z}}\mathbf{\Lambda}_{\mathbf{Z}}\mathbf{\Sigma}^{-1/2}\left(\mathbf{U}_{\mathbf{Z}}^{\top}\mathbf{U}_{\mathbf{Z}}\right)\mathbf{\Sigma}^{-1/2}\mathbf{\Lambda}_{\mathbf{Z}}\mathbf{V}_{\mathbf{Z}}^{\top}\\ &=\mathbf{V}_{\mathbf{Z}}\mathbf{\Lambda}_{\mathbf{Z}}^{2}\mathbf{\Sigma}^{-1}\mathbf{V}_{\mathbf{Z}}^{\top},\end{split}$ (7)
where $\mathbf{Z}=\left(\mathbf{U}_{\mathbf{Z}}\mathbf{\Lambda}_{\mathbf{Z}}\mathbf{V}_{\mathbf{Z}}^{\top}\right)$ denotes the singular value decomposition(SVD) of the consensus bipartite graph $\mathbf{Z}$. Therefore, we only need to compute the right singular matrix of $\mathbf{Z}\in\mathbb{R}^{m\times n}$ for further clustering task (lite $k$-means for example), and this costs $\mathcal{O}(nm^{2})$.